A large asteroid of mass 71700 kg71700 kg is at rest far away from any planets or stars. A much smaller asteroid, of mass 830 kg,830 kg, is in a circular orbit about the large asteroid at a distance of 256 m256 m as a result of their mutual gravitational attraction. What is the speed ?2v2 of the second asteroid?

?2

m/sm/s

Now suppose that the first and second asteroids carry charges of +1.68 C+1.68 C and −1.68 C,−1.68 C, respectively. How fast ?′2 would the second asteroid have to be moving in order to occupy the same circular orbit as before?

?′2=
Answer #1

a)

From Newton's second law ,gravitational force is balanced by centripetal force

GMm/r2  = mV22/r

=>V2=sqrt(GM/r) =sqrt[(6.673*10-11)(71700)/(256)]

V2=1.367*10-4 m/s

b)

From Newton's second law ,electrostatic force is balanced by centripetal force

KQ2/r2  = mV2'2/r

=>V2'=sqrt(KQ2/rm)=sqrt[(9*109)(1.68)2/(256)(830)]

V2'=345.76 m/s